OpenStudy (anonymous):

Solve for x in the proportion x -5 ------ = ----- 2x+5 x A. x=5 B. x=15 C. x=-25 D. x=-5

Parth (parthkohli):

Cross multiply.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow x^2 = -5(2x + 5) }\) \(\Large \color{MidnightBlue}{\Rightarrow x^2 = -10x - 25 }\)

Parth (parthkohli):

This makes a quadratic equation.

Parth (parthkohli):

Do you know how to solve them?

Parth (parthkohli):

Well here I give the formula. \(\Large \color{MidnightBlue}{\Rightarrow x = {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (anonymous):

Thanks for the hints, but can you just give me the answer? I'm kind of in a hurry to get the answer...

Parth (parthkohli):

Hmm if you're in a hurry and you want the answer, then I don't recommend this site to you. Use wolfram.

OpenStudy (radar):

\[x ^{2}+10x+25=0\] Same equation but in a more familiare form

Parth (parthkohli):

This is actually a perfect square, so you have only one solution.

OpenStudy (radar):

Note that it will factor, making an easy solution.

OpenStudy (anonymous):

So... what is it?

OpenStudy (radar):

(x+5)(x+5)=0 There it is factored for you.